3y2 + ky + 4
Find the value of k that makes the above trinomial factorable

Found 2 solutions by Edwin McCravy, solver91311:
Answer by Edwin McCravy(20055)

3y² + ky + 4

The only way to factor 3 is 3∙1 or 1∙3 but we can always reverse the
two binomial factors, so we might as well say any factorization is
like this:

(3y ± ?)(y ± ?)

The integers that go where the ?'s have the same sign and are 
pairs of factors that have product 4.  So they could be ±1 and ±4, ±2 and ±2, or
±4 and ±1

(3y ± 1)(y ± 4) = 3y² ± 7y + 4, so k = ±7

(3y ± 2)(y ± 2) = 3y² ± 8y + 4, so k = ±8

(3y ± 1)(y ± 4) = 3y² ± 13y + 4, so k = ±13

So there are 6 answers, 3 positive and 3 negative.

Edwin

Answer by solver91311(24713)



Presuming that you meant

There are several possibilities:













If you allow

Then one possibility is:



There are many, many others.


John

My calculator said it, I believe it, that settles it